# AI Safety in the Eyes of the Downstream Developer: A First Look at Concerns, Practices, and Challenges

### v1

@@ Abstract.
Pre-trained models (PTMs) have become a cornerstone of AI-based software, allowing for rapid integration and development with minimal training overhead. However, their adoption also introduces unique safety challenges, such as data leakage and biased outputs, that demand rigorous handling by downstream developers. While previous research has proposed taxonomies of AI safety concerns and various mitigation strategies, how downstream developers address these issues remains unexplored.
This study investigates downstream developers’ concerns, practices and perceived challenges regarding AI safety issues during AI-based software development. To achieve this, we conducted a mixed-method study, including interviews with 18 participants, a survey of 86 practitioners, and an analysis of 874 AI incidents from the AI Incident Database. Our results reveal that while developers generally demonstrate strong awareness of AI safety concerns, their practices, especially during the preparation and PTM selection phases, are often inadequate. The lack of concrete guidelines and policies leads to significant variability in the comprehensiveness of their safety approaches throughout the development lifecycle, with additional challenges such as poor documentation and knowledge gaps, further impeding effective implementation. Based on our findings, we offer suggestions for PTM developers, AI-based software developers, researchers, and policy makers to enhance the integration of AI safety measures.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06copyright: none
@@ 1. Introduction
From autonomous vehicles to powerful chatbots, modern society is adopting AI-based software (Martínez-Fernández et al., 2022) in critical domains, including education (Ciolacu et al., 2018), healthcare (Zhou et al., 2018), and transportation (Menze and Geiger, 2015).
@@
With recent advancements in AI models, the expertise and resources required to train a model from scratch have increased (Goyal, 2017; Wei et al., 2024; Patterson et al., 2021). This has led to a paradigm shift, with developers increasingly adopting pre-trained models (PTMs) from platforms such as Hugging Face (Jiang et al., 2023; Tan et al., 2022). Neglecting the safety of AI-based software has harmed individuals and broader societies (inc, [n. d.]). AI safety (Amodei et al., 2016) — avoiding harms by accidental (Amodei et al., 2016) or deliberate (Hernández-Orallo et al., 2020) mis-use — is an important factor to consider during the software development process.
@@
Many recent endeavours from the research community propose taxonomies of AI safety concerns (Amodei et al., 2016; Gao et al., 2024), evaluation methods (Wang et al., 2023), and mitigation approaches (Yi et al., [n. d.]; Roselli et al., 2019; Ji et al., 2023). Additionally, policy makers are also making efforts to provide regulations on AI safety and responsible AI design and development (eur, [n. d.]; iso, [n. d.]). Despite these efforts, how AI-based software developers perceive AI safety concerns, as well as their practices and challenges during the development phase, is not investigated. Due to the complexity of AI models and the knowledge gap between developers and model designers (Nahar et al., 2022; Kim et al., 2017), developers may approach AI safety differently and face various challenges. This is particularly true for smaller-scale company and open-source developers, where fewer resources are provided and regulations might not be enforced. This gap between AI safety frameworks and real-world practices highlights the need for an investigation into developers’ workflows, challenges, and potential shortcomings to better inform researchers and policymakers in supporting AI safety implementation.
@@
This study aims to investigate the concerns that downstream developers have about AI safety, the approaches they adopt, and the challenges they face throughout the entire AI-based software development lifecycle. To achieve this, we conducted a mixed-method study, beginning with interviews with 18 practitioners experienced in developing AI-based software, each lasting approximately one hour. We then performed a thematic analysis to derive answers to our research questions about concerns, practices, and perceived challenges when handling AI safety issues. To complement the qualitative findings, we designed and conducted a survey with 86 practitioners to triangulate our results, providing quantitative evidence on the importance of AI safety concerns, the frequency of different safety-related actions, and the extent to which developers agree on various AI safety challenges. Finally, we compared developers’ concerns against 874 user-reported AI safety issues from the AI Incident Database, identifying gaps in awareness and practices.
@@
Our results show that downstream developers’ AI safety concerns focus on data, model behaviour, and misuse and exploitation concerns. Despite developers’ generally high awareness of AI safety concerns, their actions in the early stages, particularly preparation and PTM selection, are often inadequate. Additionally, their actions are not systematic and vary significantly in comprehensiveness. The primary challenge stems from insufficient policy and regulation guidance, which is scattered across the entire development lifecycle. Other challenges, including lack of standardised benchmarks and evaluation metrics, and insufficient monitoring and maintenance methodologies, further hinder effective AI safety practices. Based on our findings, we provide concrete suggestions to AI-model developers, developers of AI-based software, researchers, and policy makers. The key contributions of our research are as follows.
@@
Significance: Software is used to perform increasingly critical functions in our societies, and it increasingly incorporates AI components to do so. Just as they have done for traditional software, software engineers must reason about the safety of the AI-based systems they develop. Although AI researchers and governments have offered guidance on “AI safety”, to date we lack the perspective of the engineers building the systems. Our study gives a first look at their concerns, practices, and challenges, setting the stage for new knowledge and improved knowledge transfer to increase the safety of the AI-based systems on which we depend.
@@ 2. Background and Related Work
Here we discuss software engineering practices for AI-based software (section 2.1), and examine the literature on AI safety (section 2.2).
@@ 2.1. Engineering Practices for AI-Based Software
These discussions underscore some of the difficulties of developing AI-based software, whether directly or via re-use. However, literature does not describe how developers actually address these concerns. In this work, we fill this gap by examining challenges and mitigation approaches adopted by PTM users through qualitative and quantitative analysis.
@@ 2.2.1. What is AI safety?
By “AI safety” we mean that an AI-based system is safe following the definition in the previous paragraph. Researchers have described many aspects of safety for AI-based systems, including discrimination (Sheng et al., 2021), misinformation, and over-reliance or unsafe use of software (Weidinger et al., 2021). Some authors also include non-accidental risks stemming from malicious use or adversarial attacks, sometimes referred to as AI security risks (Hernández-Orallo et al., 2020). With the rapid advancement and increasing real-world deployment of AI technologies, safety risks are emerging in the delivered software products. A recent systematic literature review on AI safety underscores the gaps in human value alignment, ethical governance, and safety frameworks for AI system implementation (Salhab et al., 2024). Additionally, production-grade chatbots such as ChatGPT and Gemini have raised safety concerns, including policy compliance issues (Hacker et al., 2023), susceptibility to generating or endorsing violent content (Zhuo et al., 2023), and systemic bias (Fleisig et al., 2024).
2.2.2. Realising AI Safety
@@
Researchers have proposed methods for approaching AI safety concerns during the development of AI-based software, including dataset construction methods (Wang and Levy, 2024), evaluation of AI safety risks (Wang et al., 2023), mitigation strategies (Jung et al., 2024; Chen and Liu, 2023), and monitoring strategies (Naveed et al., 2024). Additionally, governance frameworks have been proposed to address AI safety concerns. For instance, Agarwal et al. (Agarwal and Agarwal, 2024) introduced a seven-layer model to standardise fairness assessment, and Lu et al. (Lu et al., 2024) described responsible AI patterns for governance and engineering. Others have surveyed bias in AI systems across scientific disciplines (McGovern et al., 2024), and proposed governance structures such as independent audits (Falco et al., 2021) and ethics-to-practice frameworks (Shneiderman, 2020; Amodei et al., 2016). Additionally, policies like the EU AI Safety Standards (Pouget and Zuhdi, 2024) or ISO 42001 (iso, [n. d.]) aim to regulate high-risk AI systems at a higher level without concrete action suggestions (Pouget and Zuhdi, 2024).
@@
Despite ongoing efforts from AI researchers and policymakers for AI safety, little attention has been paid to understanding the engineering practices of downstream developers. Gaining insights from their perspective is crucial, as it reveals the state of practice and the challenges in addressing AI safety. This understanding helps identify and bridge gaps in knowledge (and knowledge transfer) between AI research, regulations, and real-world software engineering practices.
@@ 3. Research Questions & Methodology
Figure 1. Study design and methodology.
@@
    What are software engineers’ safety concerns in AI-based software development?
@@
To address these RQs, we conducted a mixed-method study, as illustrated in Figure 1. We began by interviewing 18 participants, as interviews are well-suited for exploring complex experiences and eliciting rich, detailed insights from downstream developers (Jiang et al., 2023). This was followed by a survey with 86 participants to validate and quantify the findings (Aniche et al., 2018), and an analysis of 874 incident cases from the AI Incident Database to triangulate our results against real-world AI safety failures (Anandayuvaraj et al., 2024).
In the interview, we did not explicitly define “AI Safety” for the participants. Given the sufficiently abstract concept, we instead asked the concerns they have during development to avoid subjectivity. We further calibrate if the participants’ answers go far beyond general AI safety scopes (Hernández-Orallo et al., 2020), such as carbon emission or super-intelligence replacing humans (“singularity”). In the survey instrument, we asked more concretely about AI safety concerns derived from the interview results.
@@ 3.1. Method 1: Interviews
Instrument Design. We designed a semi-structured interview protocol exploring the three research questions. Following the AI-based software development process outlined by Amershi et al. (Amershi et al., 2019), we structured the interview into sections covering background understanding, model selection, model development and testing, model deployment and maintenance. For each stage, we prepared questions asking participants’ development experience, concerns, practices, and challenges. We conducted three pilot interviews with AI-based software developers and then revised our interview protocol based on the results. The final version of our interview took approximately one hour. Table 1 shows sample questions.
Table 1. Summary of interview and survey instruments. (See complete data in replication package)
@@
Recruitment. To access downstream developers actively engaged in AI-based software development, we recruited participants from the Hugging Face ecosystem and our professional network, targeting those with AI-based software development experience. Following the design of a previous study to recruit participants from Hugging Face (Jiang et al., 2024a), we extracted all PRO users and members of organisations with over 50 users, retaining only those labelled as company or non-profit, excluding labels such as community and university. Since Hugging Face profiles sometimes link to personal websites, we contacted users via publicly available emails on their personal websites to comply with site policies (git, [n. d.]). This process yielded 98 PRO users and 783 organisational users, totalling 868 unique users after deduplication.
@@
We first contacted PRO users, followed by organisational users, prioritising those from larger organisations based on the number of members. Recruitment and analysis occurred concurrently. In total, we contacted 667 users on Hugging Face. 17 users scheduled interviews with us, of which 13 attended. This resulted in a response rate of 2.55% similar to other software engineering studies (Lethbridge et al., 2005; Singer et al., 2014). To complement this sample, we invited five industry AI-based software developers from our personal network, totalling 18 participants. Table 2 shows the demographics of our interview participants. Each participant was compensated \$20 USD for their time.
Table 2. Interview participant demographics.
ID DL Exp. SE Exp. # AI software Degree Org. Size Domain
P1 1 15 1-5 Bachelor Open-source NLP
P2 6 6 5-10 PhD Small CV
P3 2 4 1-5 Bachelor Open-source NLP
P4 1 2 1-5 PhD Large NLP
P5 2 7 1-5 PhD Open-source CV
P6 12 25 1-5 PhD Small NLP
P7 4 2 10-25 PhD Large NLP
P8 2 2 25+ Master Large CV
P9 6 6 1-5 PhD Large CV, NLP
P10 7 4 5-10 PhD Large CV
P11 4 3 25+ Master Medium CV, NLP
P12 4 4 25+ Master Large CV, NLP
P13 3 2 5-10 Bachelor Large NLP
P14 1 2 5-10 PhD Small NLP
P15 3 3 1-5 Bachelor Medium NLP
P16 1 2 25+ Master Small NLP
P17 4 4 25+ PhD Small NLP
P18 7 7 10-25 PhD Medium NLP
Figure 2. Illustration of our coding process.
@@
Analysis. Following suggestions in software engineering interview studies (Hove and Anda, 2005), the first two sessions were conducted by the first author and a senior researcher, who is also co-author of this study. They discussed the interview flow, after which the first author conducted the remaining sessions. We transcribed the interview recordings using Microsoft 365’s transcription tool, verified manually, and organised in NVivo for analysis. Then we adopted suggested steps from thematic analysis (Braun and Clarke, 2006; Cruzes and Dyba, 2011) for the qualitative analysis stage of the interview transcripts. Two authors, including the first author and the author who participated in two interview sessions and closely monitored the interview transcripts, first read through two interview transcripts, summarising the content as key points related to AI safety concerns, practices, and challenges that corresponded to the three research questions, as suggested to initiate the open coding (Hoda et al., 2012). They then proceeded to assign base level of codes to the extracted key points. Subsequently, we developed higher level abstractions of subthemes by grouping relevant codes together and then the highest level abstractions of themes. Consensus was reached before the first author continued to apply this process to the rest of the transcripts. During the coding process, we constantly compared the emerging codes within and across interview transcript key points, with regular discussions held to verify the code assignment. This iterative process of code development led to continuous adjustment of the codes. Grouping of the codes into themes was conducted concurrently, with each subtheme and theme discussed and adjusted as well. Finally, we derived 29, 100, and 55 codes for the three RQs, respectively. Three authors independently reviewed the coding process and the synthesised subthemes and themes for each research question. The results received agreement among the entire group. Figure 2 displays a sample of the coding process. To determine whether additional participants are needed, we measured the saturation. Following Guest et al.’s (Guest et al., 2006) recommendation, we measured saturation by tracking the cumulative appearance of subthemes in each interview. Saturation in all RQs was achieved after interview 13.

### v3

@@ Abstract.
With the advancement of AI models, more software systems are adopting AI as a component to facilitate automation. Pre-trained models (PTMs) have become a cornerstone of AI-based software, allowing for rapid integration and development with lower training cost. However, their adoption also introduces unique safety challenges, such as data leakage and biased outputs, that may require careful handling by downstream developers. While previous research has proposed taxonomies of AI safety concerns and various mitigation strategies, how downstream developers address these issues during the development of AI-based software when reusing PTMs remains unexplored. Understanding downstream developers’ perspectives is essential because they directly influence how AI safety concerns translate into practice, such as determining whether risks like data leakage or model bias are recognised, mitigated, or inadvertently overlooked in real-world deployments.
This study investigates downstream developers’ concerns, practices and perceived challenges regarding AI safety issues during the development of AI-based software. To achieve this, we conducted a mixed-method study, including interviews with 16 participants, a survey of 86 practitioners, and an analysis of 874 AI incidents from the AI Incident Database. Our results reveal that while developers generally demonstrate strong awareness of AI safety concerns, their practices, especially during the preparation and PTM selection phases, are often inadequate. The lack of concrete guidelines and policies leads to significant variability in the comprehensiveness of their safety approaches throughout the development lifecycle, with additional challenges such as poor documentation and knowledge gaps, further impeding effective implementation. Based on our findings, we offer suggestions for PTM developers, developers of AI-based software, researchers, and policy makers to enhance the integration of AI safety measures.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
From autonomous vehicles to powerful chatbots, software engineers are adopting AI models to build AI-based software systems (Martínez-Fernández et al., 2022) in critical domains, including education (Ciolacu et al., 2018), healthcare (Zhou et al., 2018), and transportation (Menze and Geiger, 2015). With recent advancements in AI models, the expertise and resources required to train a model from scratch have increased (Goyal, 2017; Wei et al., 2024; Patterson et al., 2021). This has led to a paradigm shift, with developers increasingly adopting pre-trained models (PTMs) — models that have been trained on large datasets for general tasks and can be adapted to specific applications — from platforms such as Hugging Face (Jiang et al., 2023; Tan et al., 2022). Neglecting the safety of AI-based software has harmed individuals and broader societies (inc, [n. d.]). AI safety (Amodei et al., 2016) — avoiding harms by accidental (Amodei et al., 2016) or deliberate (Hernández-Orallo et al., 2020) mis-use — is an important factor to consider during the software development process. The way downstream developers handle AI safety throughout the development process, including data collection, fine-tuning, evaluation, deployment, and monitoring (Amershi et al., 2019), directly shapes the risks and reliability of the final software.
@@
Many recent endeavours from the research community propose taxonomies of AI safety concerns (Amodei et al., 2016; Gao et al., 2024), evaluation methods (Wang et al., 2023), and mitigation approaches (Yi et al., [n. d.]; Roselli et al., 2019; Ji et al., 2023). Additionally, policy makers are also making efforts to provide regulations on AI safety and responsible AI design and development (eur, [n. d.]; iso, [n. d.]). Despite these efforts, how developers of AI-based software perceive AI safety concerns, as well as their practices and challenges during the development phase in the context of reusing PTMs, is not investigated. Due to the complexity of AI models and the knowledge gap between developers and model designers (Nahar et al., 2022; Kim et al., 2017), developers may approach AI safety differently and face various challenges. This is particularly true for smaller-scale company and open-source developers, where fewer resources are provided and regulations might not be enforced. This motivates the need to examine whether and how current AI safety frameworks align with or differ from the real-world development practices, as well as potential shortcomings to better inform researchers and policymakers in supporting AI safety implementation.
@@
This study aims to investigate the concerns that downstream developers have about AI safety, the approaches they adopt, and the challenges they face throughout the entire development lifecycle of AI-based software. To achieve this, we conducted a mixed-method study, beginning with interviews with 16 practitioners experienced in developing AI-based software, each lasting approximately one hour. We then performed a thematic analysis to derive answers to our research questions about concerns, practices, and perceived challenges when handling AI safety issues. To complement the qualitative findings, we designed and conducted a survey with 86 practitioners to triangulate our results, providing quantitative evidence on the importance of AI safety concerns, the frequency of different safety-related actions, and the extent to which developers agree on various AI safety challenges. Finally, we compared developers’ concerns against 874 user-reported AI safety issues from the AI Incident Database, identifying gaps in awareness and practices.
@@
Our results show that downstream developers’ AI safety concerns focus on data, model behaviour, and misuse and exploitation concerns. Additionally, their awareness of these AI safety concerns is high in general. The practices identified by downstream developers include deeper understanding and assessment for safety concerns, regular monitoring and assessment for safety concerns, implementing technical safety guards, maintaining documentation and expert consultation and relying on external parties and resources. However, these actions are not systematic and vary significantly in comprehensiveness. In fact, despite developers’ generally high awareness of AI safety concerns, their actions in the early stages, particularly preparation and PTM selection, are often inadequate. The primary challenge stems from insufficient policy and regulation guidance, which is scattered across the entire development lifecycle. Other challenges, including lack of standardised benchmarks and evaluation metrics, and insufficient monitoring and maintenance methodologies, further hinder effective AI safety practices. Based on our findings, we provide concrete suggestions to AI-model developers, developers of AI-based software, researchers, and policy makers. To address the knowledge transfer gap, we suggest addressing the knowledge transfer gap by urging AI model developers to improve documentation and by advocating for a practical AI safety checklist developed by researchers and policymakers, to guide downstream developers.
In summary, the key contributions of our research are as follows.
@@
Significance to Software Engineering: Software is used to perform increasingly critical functions in our societies, and it increasingly incorporates AI components to do so. Just as software engineers have reasoned about the safety of the traditional software systems they developed, they must now reason about the safety of their AI-based systems. Although AI researchers and governments have offered guidance on “AI safety”, to date we lack the perspective of the engineers building the systems. Our study gives a first look at their concerns, practices, and challenges, setting the stage for new knowledge and improved knowledge transfer to increase the safety of the AI-based systems on which we depend. We identify the knowledge transfer gap affecting downstream developers. For PTM developers, improved documentation is needed to enhance transparency and usability. For policymakers, we recommend the development of evolving AI safety checklists, accompanied by actionable guidance linked to the most effective and up-to-date mitigation strategies. For researchers, we suggest focusing on bridging this knowledge transfer gap by investigating the specific challenges faced by downstream users, identifying the most useful types of knowledge, and exploring effective modes of knowledge dissemination.
@@ 2. Background and Related Work
Here we discuss software engineering practices for AI-based software (section 2.1), examine the literature on AI safety (section 2.2), finally compare it with the broader concept of AI ethics and discuss relevant studies on developers’ perspectives to help scope and distinguish our study (section 2.3).
@@ 2.1. Engineering Practices for AI-Based Software
These discussions underscore some of the difficulties of developing AI-based software, whether directly or via re-use, with AI safety concerns emerging as a key issue. However, literature does not describe how developers actually address the AI safety concerns. In this work, we fill this gap by examining challenges and mitigation approaches adopted by PTM users through qualitative and quantitative analysis.
@@ 2.2. AI Safety Concerns and Mitigations
Researchers have described many aspects of safety for AI-based systems, including discrimination (Sheng et al., 2021), misinformation (Zhang et al., 2023), and over-reliance or unsafe use of software (Weidinger et al., 2021). Some authors also include non-accidental risks stemming from malicious use or adversarial attacks, sometimes referred to as AI security risks (Hernández-Orallo et al., 2020). With the rapid advancement and increasing real-world deployment of AI technologies, safety risks are emerging in the delivered software products. A recent systematic literature review on AI safety underscores the gaps in human value alignment, ethical governance, and safety frameworks for AI system implementation (Salhab et al., 2024). Additionally, production-grade chatbots such as ChatGPT and Gemini have raised safety concerns, including policy compliance issues (Hacker et al., 2023), susceptibility to generating or endorsing violent content (Zhuo et al., 2023), and systemic bias (Fleisig et al., 2024).
Meanwhile, the engineering practices for ensuring safety in traditional software become insufficient when applied to AI-based software. For example, comprehensive tests for validating and verifying a system meets the safety requirements is a standard practice in conventional software. However, AI systems often work in a more open-world setting, encountering diverse and unanticipated inputs that cannot be exhaustively enumerated in advance. Moreover, the model behaviour could drift or degrade when data distribution change over time. Therefore, dedicated practices tailored to the unique characteristic of AI-based systems are necessary to ensure safety.
@@
Researchers have proposed methods for approaching AI safety concerns during the development of AI-based software, including dataset construction methods (Wang and Levy, 2024), evaluation of AI safety risks (Wang et al., 2023), mitigation strategies (Jung et al., 2024; Chen and Liu, 2023), and monitoring strategies (Naveed et al., 2024). Additionally, governance frameworks have been proposed to address AI safety concerns. For instance, Agarwal et al. (Agarwal and Agarwal, 2024) introduced a seven-layer model to standardise fairness assessment, and Lu et al. (Lu et al., 2024) described responsible AI patterns for governance and engineering. Others have surveyed bias in AI systems across scientific disciplines (McGovern et al., 2024), and proposed governance structures such as independent audits (Falco et al., 2021) and ethics-to-practice frameworks (Shneiderman, 2020; Amodei et al., 2016). Furthermore, policies such as the EU AI Safety Standards (Pouget and Zuhdi, 2024) or ISO 42001 (iso, [n. d.]) aim to regulate high-risk AI systems. ISO/IEC 42001 (iso, [n. d.]) is an international standard for AI systems, offering a framework for organisations address risks related to transparency, robustness, and accountability. Similarly, ISO/IEC 15408 (EAL, [n. d.]) defines evaluation assurance levels used to measure the confidence that system’s principal features are reliably implemented. While primarily used for general software assurance, these levels, from EAL1 (functionally tested) to EAL7 (formally verified design and tested) illustrate how multi-tiered assessment framework can be adopted to evaluate AI systems. However, such policies often remain high-level and do not provide concrete action suggestions (Pouget and Zuhdi, 2024).
@@
Despite ongoing efforts from AI researchers and policymakers for AI safety, little attention has been paid to understanding the safety-related engineering practices of downstream developers. Gaining insights from their perspective is crucial, as it reveals the state of practice and the challenges in addressing AI safety. This understanding helps identify and bridge gaps in knowledge (and knowledge transfer) between AI research, regulations, and real-world software engineering practices.
2.3. The Scope of Our Study
While AI safety is often considered part of AI ethics, the two are not interchangeable. AI safety typically focuses on preventing unintended harm through engineering practices—such as ensuring reliability, robustness, and predictable failure—whereas AI ethics includes broader societal values like fairness, accountability, and transparency. AI ethics is broad in scope, defined by Siau & Wang (Siau and Wang, 2020) as “the principles of developing AI to interact with other AIs and humans ethically and function ethically in society”. Meanwhile, governmental regulations such as those in Australia (AUS, [n. d.]) and the EU (EU-, [n. d.]) often frame AI ethics broadly. These frameworks encompass human-centred values; human, social, and environmental well-being; transparency; and explainability, among other considerations, within which AI safety constitutes one component.
The scope of our study is AI safety, which is a relevant concept usually identified under AI ethics, as discussed above. Within the broader scope of AI ethics, previous studies have examined developers’ contexts to understand their perspectives. Pant et al. (Pant et al., 2024) conducted a closed-ended survey with 100 general AI practitioners to examine their awareness of AI ethics, and the challenges they face when incorporating ethics into AI development. They identified three high-level categories of challenges: (1) general challenges, (2) technology-related challenges, and (3) human-related challenges. Ali et al. (Ali et al., 2023) interviewed 25 AI ethics entrepreneurs to investigate the practical difficulties of implementing AI ethics within technology companies. Their findings highlight organisational issues such as unstable team structures, the decoupling of practices from ethical principles, and the low prioritisation of ethical concerns. Wong (Wong, 2021) also investigated user experience (UX) professionals’ attempts to reshape the values embodied and promoted by their companies.
Different from the previously mentioned works, our study focuses on AI safety, which is a more specific domain than general AI ethics. Rather than targeting general AI practitioners, we investigate the dominant pattern of PTM reuse in the development of AI-based software (Banyongrakkul et al., 2025). Accordingly, we recruited participants who help embed intelligent components into the systems they ship, where those components were derived in some way from PTMs. The integration of PTMs into a software system involves multiple development steps. This study offers a focused perspective on downstream developers engaged in PTM reuse for the development of AI-based software, particularly in the context of the entire development lifecycle (SDLC).
@@ 3. Research Questions & Methodology
Figure 1. Study design and methodology
@@
    What are software engineers’ safety concerns in the development of AI-based software?
@@
To address these RQs, we conducted a mixed-method study, as illustrated in Figure 1. We began by interviewing 16 participants, as interviews are well-suited for exploring complex experiences and eliciting rich, detailed insights from downstream developers (Jiang et al., 2023). This was followed by a survey with 86 participants to validate and quantify the findings (Aniche et al., 2018), and an analysis of 874 incident cases from the AI Incident Database to triangulate our results against real-world AI safety failures (Anandayuvaraj et al., 2024).
In the interviews, to observe the downstream developers’ concerns, we let participants articulate their concerns without imposing a specific definition to avoid bias, as suggested when conducting interview studies (Charmaz, 2006). Internally, we borrowed the AI safety definition by Hernandez-Orallo et al. (Hernández-Orallo et al., 2020), which includes “both risks from AI systems for which the source of risk is accidental, ranging from unpredictable systems to negligent use, and non-accidental risks such as those stemming from malicious use or adversarial attacks”. If needed, during the interview, we steered the conversation into the direction of this AI safety definition. For example, when the participants’ answers went far beyond general AI safety scopes (Hernández-Orallo et al., 2020), such as carbon emissions or super-intelligence replacing humans (“singularity”), we calibrated the response. In the survey, giving a specific definition was not necessary since all questions in the survey were directly derived from findings from the interviews.
@@ 3.1. Method 1: Interviews
Instrument Design. We designed a semi-structured interview protocol exploring the three research questions. Following the development process of AI-based software outlined by Amershi et al. (Amershi et al., 2019), we structured the interview into sections covering background understanding, model selection, model development and testing, model deployment and maintenance. For each stage, we prepared questions asking participants’ development experience, concerns, practices, and challenges. We conducted three pilot interviews with developers of AI-based software and then revised our interview protocol based on the results. The final version of our interview took approximately one hour. Table 1 shows sample questions.
Table 1. Summary of interview and survey instruments. (See complete data in replication package section 8)
@@
Recruitment. To access downstream developers actively engaged in the development of AI-based software, we recruited participants from the Hugging Face ecosystem and our professional network, targeting those with development experience in AI-based software. Following the design of a previous study to recruit participants from Hugging Face (Jiang et al., 2024a), we extracted all PRO users and members of organisations with over 50 users, retaining only those labelled as company or non-profit, excluding labels such as community and university. Since Hugging Face profiles sometimes link to personal websites, we contacted users via publicly available emails on their personal websites to comply with site policies (git, [n. d.]). This process yielded 98 PRO users and 783 organisational users, totalling 868 unique users after deduplication.
@@
We first contacted PRO users, followed by organisational users, prioritising those from larger organisations based on the number of members. Recruitment and analysis occurred concurrently. In total, we contacted 667 users on Hugging Face. 17 users scheduled interviews with us, of which 13 attended. This resulted in a response rate of 2.55% similar to other software engineering studies (Lethbridge et al., 2005; Singer et al., 2014). During the interviews, 2 participants did not provide useful responses, and we discarded them from the analysis. To complement this sample, we invited five industry developers of AI-based software from our personal network, totalling 18 participants, with 16 being analysed. Table 2 shows the demographics of our interview participants, in which we keep the ID for all 18 participants, but removed P3 and P5 from the analysis. Each participant was compensated \$20 USD for their time. Additionally, interviews were conducted and analysed by researchers with technical and qualitative training, including experience in AI-based software development and AI safety research.
Table 2. Interview participant demographics. The unit of DL and SE Exp. are in years. See section 3.4 for safety-critical definition. P3 and P5 were excluded from the analysis.
ID DL Exp. SE Exp. # AI software Degree Org. Size Domain Safety-Critical
P1 1 15 1-5 Bachelor Open-source NLP No
P2 6 6 5-10 PhD Small CV No
P3 2 4 1-5 Bachelor Open-source NLP No
P4 1 2 1-5 PhD Large NLP No
P5 2 7 1-5 PhD Open-source CV No
P6 12 25 1-5 PhD Small NLP No
P7 4 2 10-25 PhD Large NLP Yes
P8 2 2 25+ Master Large CV No
P9 6 6 1-5 PhD Large CV, NLP No
P10 7 4 5-10 PhD Large CV No
P11 4 3 25+ Master Medium CV, NLP No
P12 4 4 25+ Master Large CV, NLP No
P13 3 2 5-10 Bachelor Large NLP No
P14 1 2 5-10 PhD Small NLP Yes
P15 3 3 1-5 Bachelor Medium NLP No
P16 1 2 25+ Master Small NLP No
P17 4 4 25+ PhD Small NLP No
P18 7 7 10-25 PhD Medium NLP No
Figure 2. Illustration of our coding process
@@
Analysis. Following suggestions in software engineering interview studies (Hove and Anda, 2005), the first two sessions were conducted by the first author and a senior researcher, who is also co-author of this study. They discussed the interview flow, after which the first author conducted the remaining sessions. We transcribed the interview recordings using Microsoft 365’s transcription tool, verified manually, and organised in NVivo for analysis. Then we adopted suggested steps from thematic analysis (Braun and Clarke, 2006; Cruzes and Dyba, 2011) for the qualitative analysis stage of the interview transcripts. Two authors, including the first author and the author who participated in two interview sessions and closely monitored the interview transcripts, first read through two interview transcripts, summarising the content as key points related to AI safety concerns, practices, and challenges that corresponded to the three research questions, as suggested to initiate the open coding (Hoda et al., 2012). They then proceeded to assign base level of codes to the extracted key points. Subsequently, we developed higher level abstractions of subthemes by grouping relevant codes together and then the highest level abstractions of themes. Consensus was reached before the first author continued to apply this process to the rest of the transcripts. During the coding process, we constantly compared the emerging codes within and across interview transcript key points, with regular discussions held to verify the code assignment. This iterative process of code development led to continuous adjustment of the codes. Grouping of the codes into themes was conducted concurrently, with each subtheme and theme discussed and adjusted as well. Finally, we derived 29, 100, and 55 codes for the three RQs, respectively. Three authors independently reviewed the coding process and the synthesised subthemes and themes for each research question. The results received agreement among the entire group. Figure 2 displays a sample of the coding process. To determine whether additional participants are needed, we measured the saturation. Following Guest et al.’s (Guest et al., 2006) recommendation, we measured saturation by tracking the cumulative appearance of subthemes in each interview. Saturation in all RQs was achieved after interview 13. Figure 3 shows the saturation curve for the number of unique subthemes.
Figure 3. Saturation Curve for Interview Participants